OpenStudy (anonymous):

Given the x and y intercept, write the equation of the line in the form ax+by=c x=6 y= -2 my answer is x-12y=12 am i right?

OpenStudy (kainui):

I don't think that's quite right. Show me your steps and I'll see if I can help you find where you went wrong.

OpenStudy (kainui):

For example, if that's your y-intercept y=-2, then x=0 at that point. If we plug this into your equation we get: x-12y=12 (0)-12(-2)=12 24=12 Well that's not true! So you're slightly off somewhere.

OpenStudy (anonymous):

@Pandaaa: Hi ! Did you get it ?

OpenStudy (anonymous):

Need my help ?

OpenStudy (anonymous):

Yes please? :) can you show me how to do it using the numbers i used?

OpenStudy (anonymous):

Sure ;) Just look : This is an example : If We have x=1 and y=2 : x+y = ? have : 1 + 2 = 3 Or : X^2+3y=? 1^2+3*2=1+9=10 Got it ?;)

OpenStudy (anonymous):

I kinda got confused. Is this how to right the form ax+by=c ?

OpenStudy (anonymous):

Step by step friend ;)